# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v3

@@ 5.3.2. Model Training
We fine-tune two instruction-tuned backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.4. Query Clustering Task
After training, we apply the fine-tuned models to encode all queries from the clustering benchmarks. Pairwise distances between queries are computed using cosine similarity in the embedding space. Following (Tang et al., 2022), we evaluate clustering quality using the BetaCV (Mohammed and Wagner, 2014) metric, which measures intra-cluster compactness relative to inter-cluster distance. Specifically, a lower BetaCV score indicates better clustering performance, as it reflects that queries within the same cluster are more tightly grouped, while those from different clusters are more widely separated.

### v4

@@ 5.3.2. Model Training
We fine-tune two backbone models, Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, on the constructed dataset of ¡NL, SQL¿ pairs. Both models are pretrained for code and SQL reasoning, making them suitable for Text-to-SQL adaptation. Training is performed using the AdamW optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and $\epsilon=1\times 10^{-8}$. We use a peak learning rate of $3\times 10^{-6}$, linearly warmed up over the first 5% of training steps, followed by cosine learning rate decay. Each model is trained for 3 epochs with a batch size of 256 and a maximum sequence length of 4096 tokens. To stabilize training and prevent overfitting, we apply gradient clipping with a threshold of 1.0 and set the weight decay to 0.01. All other settings follow the default configuration of LLaMA-Factory (Zheng et al., 2024).
@@ 5.4. Query Clustering Task
After training, we apply the fine-tuned models to encode all queries from the clustering benchmarks. Pairwise distances between queries are computed using cosine similarity in the embedding space. We evaluate clustering quality using the BetaCV (Mohammed and Wagner, 2014) metric, which measures intra-cluster compactness relative to inter-cluster distance. Specifically, a lower BetaCV score indicates better clustering performance, as it reflects that queries within the same cluster are more tightly grouped, while those from different clusters are more widely separated.